An ideal gas at 0.80 atmospheres and 87°C occupies 0.450 liter. How many moles are in the sample? (R = 0.0821 liter-atm/mole – K)

n = ____ mol


1
Expert's answer
2021-09-12T18:59:26-0400

Remembering that P1V1/T1 = P2V2/T2 we have that V2 = (P1V1T2)/(T1P2), from which we can calculate the volume at 1 atm and 0 oC (273 K)

P1V1T2 = (.452)(.62)(273) = 76.5

T1P2 = (360)(1) = 360

V2 = 76.5/360 = 0.212 liters

n = PV / RT = (.452)0.62/(0.00820)(360) = 0,0094

Alternatively remembering that the volume of 1 mole at standard conditions is 22.41 liters

n = 0.212/22.41 = 0,0094